OpenStudy (anonymous):

ILL GIVE YOU A MEDAL!!! If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 6x − y = 1 4x − 3y = −11

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@bdude999

OpenStudy (anonymous):

@mathacomplice

OpenStudy (texaschic101):

6x - y = 1 6x - 1 = y now sub 6x - 1 in for y in the 2nd equation 4x - 3y = -11 4x - 3(6x - 1) = -11 <=== I think that is what you need

OpenStudy (anonymous):

Thank you!!! @texaschic101